OpenStudy (anonymous):

Find the area and the circumference of a circle with radius . Use the value for , and do not round your answers. Be sure to include the correct units in your answers.

OpenStudy (anonymous):

Find the area and the circumference of a circle with radius 9yd . Use the value for 3.14pie , and do not round your answers. Be sure to include the correct units in your answers.

terenzreignz (terenzreignz):

You know the formula for circumference?

OpenStudy (anonymous):

i forgot

terenzreignz (terenzreignz):

It's \[\Large 2\times \pi \times r\] Where is the radius.

terenzreignz (terenzreignz):

Where *r is the radius, sorry.

OpenStudy (anonymous):

9yd

OpenStudy (anonymous):

2 x 3.14 x 9=56.52 which the circumference or

terenzreignz (terenzreignz):

That's a good enough circumference. Now, to get the area, use this formula: \[\Large A = \pi \times r\times r\]

OpenStudy (anonymous):

alright thx

terenzreignz (terenzreignz):

No problem.